# Untangling Lariats: Subgradient Following of Variationally Penalized Objectives

### v1

@@ Abstract
We describe a novel subgradient following apparatus for calculating the optimum of convex problems with variational penalties. In this setting, we receive a sequence $y_{i},\ldots,y_{n}$ and seek a smooth sequence $x_{1},\ldots,x_{n}$. The smooth sequence attains the minimum Bregman divergence to an input sequence with additive variational penalties in the general form of $\sum_{i}g_{i}(x_{i+1}-x_{i})$. We derive, as special cases of our apparatus, known algorithms for the fused lasso and isotonic regression. Our approach also facilitates new variational penalties such as non-smooth barrier functions. We next derive and analyze multivariate problems in which ${\mathbf{x}}_{i},\mathbf{y}_{i}\in\mathbb{R}^{d}$ and variational penalties that depend on $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$. The norms we consider are $\ell_{2}$ and $\ell_{\infty}$ which promote group sparsity. Last but not least, we derive a lattice-based subgradient following for variational penalties characterized through the output of arbitrary convolutional filters. This paradigm yields efficient solvers for problems in which sparse high-order discrete derivatives such as acceleration and jerk are desirable.
@@ 1 Sub-Introduction
Rather than descending into a wordy introduction conceived by ESL writers which was then passed through a spell checker and further beautified by an extra-fine-tuned language model, let us delve right into a simple gateway for this paper by examining the so called fused lasso [16]. This problem is concerned with the following optimization task,
$\min_{{\mathbf{x}}\in\mathbb{R}^{n}}\frac{1}{2}\|{\mathbf{x}}-\mathbf{y}\|^{2}+\sum_{i=1}^{n}\lambda_{i}|x_{i+1}-x_{i}|~{}~{}.$ (1)
The first part of the paper is concerned with an abstract unified treatment that generalizes the setting of the fused lasso to optimization problems of the form,
$\min_{{\mathbf{x}}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}g_{i}(x_{i}-x_{i+1})~{}~{}.$
Here, both $h_{i}:\mathbb{R}\to\mathbb{R}_{+}$ and $g_{i}:\mathbb{R}\to\mathbb{R}_{+}$ are convex in their single argument. In addition to retrieving, as special cases, both the fused-lasso and isotonic regression, our algorithm apparatus can be used with non-smooth variational penalties for $g_{i}$. To illustrate its power we derive an entirely new variant with barrier functions serving as the skeleton for variational penalty. We fondly refer to this refreshingly new variant as the Untangled Lariat (if Rob could so can we).
We explored uncharted water in which each element of the input sequence is a vector and thus need to find an sequence ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}$ where ${\mathbf{x}}_{i}\in\mathbb{R}^{d}$ for all $i\in[n]$. In order to obtain variational sparsity in the multivariate case we replace the absolute value $|x_{i+1}-x_{i}|$ with norm penalties over the differences $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$. As a starting point we derive an aesthetically succinct closed form for the squared $2$-norm, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|^{2}$. We use the closed form solution to build surrogate functions for the $2$-norm itself, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$, and the infinity-norm, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}$ yielding an efficient iterative solver for the problem,
@@
For brevity of the derivation we focus on the case where $h_{i}({\mathbf{x}}_{i})=\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}$. To the best of our minute yet tangible knowledge these settings and their analyses are new and provide new data-science and machine learning tools for multivariate temporal problems with variational sparsity.
In the last part we incorporate a common construction from signal processing and penalize the output of the convolution filter applied to the input sequence $y_{1},\ldots,y_{n}$. We provide efficient apparatus for finding the optimum of,
$\min_{\mathbf{x}}\sum_{i}h_{i}(x_{i})+\sum_{i}\big{|}(\boldsymbol{\tau}\ast{\mathbf{x}})_{i+1}\big{|}~{},$
@@
where $\boldsymbol{\tau}\in\mathbb{R}^{k}$ is an arbitrary filter of $k-1$ free parameters. We introduce a subgradient following mechanism over a $k$-dimensional lattice requires $\mathcal{O}(n^{k})$ to construct in the worst case. In applications where $\mathbf{y}$ represents a temporal sequence, the variational penalty $x_{i+1}-x_{i}$ can be viewed as promoting smoothed output with regions of zero discrete velocity. Our general framework can be used as a smoothing apparatus for specific problems in which we seek regions of zero acceleration, zero jerk, and other forms of finite differences.
Before diving in, we do want to pay some homage to previous work and make connections to our framework. The core of related research preceding this paper can be traced back to the work on the fused lasso [16]. The setting of the fused lasso has been revisited numerous times and found various venues for applications, see for instance [7, 8], though any attempt to be exhaustive here is going to injustice to the plethora of applications that employ the fused lasso. Several authors [4, 11, 12] described and analyzed a linear time algorithm for solving the fused lasso problem. In addition, alternative algorithms such as the ones described in [1, 3] do not guarantee linear time yet demonstrate excellent empirical results. The most relevant paper to our work is [12] where the authors follow the derivation in [11] to generalize the algorithm to non-smooth unary functions and to trees. Their procedure is an instance of message passing algorithm [6]. The paper [12] also considers different unary functions subject to a pairwise term. The term includes the fused lasso penalty as a special case. The lens of the work is by nature probabilistic and does not discuss general pairwise penalties. We encapsulate and generalize the algorithmic work above by providing a new subgradient following algorithm for general convex unary functions and pairwise terms.
Our approach is morally different than the above work by employing the Fenchel dual of the primal problem, a view that was advocated by Shalev-Shwartz [15] in the context of online learning. This view also facilitates a simple yet nonetheless effective generalization to Bregman divergences [2] as the objective at no additional computation cost and, as mentioned above, is equally applicable to non-smooth variational penalties. We conclude with a short discussion on extensions and generalizations of this work such as incorporating the framework from [12] for general acyclic graphs to our setting. While left for future research, we believe that the approximation for general graphs of [13], that applies 1D fused lasso solver over induced chains can be applied as well. We provide in an appendix Python code that implements subgradient following procedures for many of the settings discussed in this paper.
@@ 2 Abstract Subgradient Following Algorithm
Throughout the $h_{i}$ and $g_{i}$ are convex and lower semi-continuous. To remind the reader, the Fenchel dual of a function $f$ is defined as
$f^{*}(\alpha)=\sup_{x\in\mathrm{dom}(f)}\big{\{}\alpha x-f(x)\big{\}}~{}.$
The domain of the functions we consider is closed, compact, and convex and thus we can replace the supremum with a maximum. We next apply Fenchel-Young inequality to each $g_{i}$,
$g_{i}(x_{i}-x_{i-1})+g_{i}^{*}(\alpha)\geq\alpha(x_{i}-x_{i-1})~{},$
to bound $U$ from below,
$U$$\geq$ $\min_{{\mathbf{x}}}\,\max_{\boldsymbol{\alpha}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}\left[\alpha_{i}(x_{i}-x_{i+1})-g_{i}^{*}(\alpha_{i})\right]$
$=$ $\min_{x_{1}}\,\max_{\alpha_{1}}\,\min_{x_{2}}\ldots\max_{\alpha_{n-1}}\,\min_{x_{n}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}\left[\alpha_{i}(x_{i}-x_{i+1})-g_{i}^{*}(\alpha_{i})\right]$
$=$ $\max_{\boldsymbol{\alpha}}\,\min_{{\mathbf{x}}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}\left[\alpha_{i}(x_{i}-x_{i+1})-g_{i}^{*}(\alpha_{i})\right]~{},$
where we used successive application of the minimax theorem in the two equalities above. We confine ourselves to settings for which strong duality holds thus,
$U$ $=$ $\min_{x_{1}}\,\max_{\alpha_{1}}\,\min_{x_{2}}\ldots\max_{\alpha_{n-1}}\,\min_{x_{n}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}\left[\alpha_{i}(x_{i}-x_{i+1})-g_{i}^{*}(\alpha_{i})\right]$
@@
$f_{i}(x_{i})$$\equiv$ $\max_{\alpha_{i}}\,\ldots\min_{x_{n}}\>\sum_{j=i}^{n-1}\left[h_{j+1}(x_{j+1})+\alpha_{j}(x_{j}-x_{j+1})-g_{j}^{*}(\alpha_{j})\right]~{}.$ (3)
For boundary conditions we let $\forall x:f_{n}(x)=0$. In addition, we introduce the auxiliary function $\tilde{f}_{i}(x_{i})=f_{i}(x_{i})+h_{i}(x_{i})$. Using these definitions we introduce a summation-free form for $f_{i}$,
@@
$=$ $\max_{\alpha_{i}}\>\alpha_{i}x_{i}-g_{i}^{*}(\alpha_{i})+\min_{x_{i+1}}\left\{\tilde{f}_{i+1}(x_{i+1})-\alpha_{i}x_{i+1}\right\}$
$=$ $\max_{\alpha_{i}}\>\alpha_{i}x_{i}-g_{i}^{*}(\alpha_{i})-\tilde{f}_{i+1}^{*}(\alpha_{i})~{}.$
Note that to find $f_{i}(x_{i})$ we need to obtain the minimum of $\tilde{f}_{i+1}(x_{i+1})-\alpha_{i}x_{i+1}$ with respect to $x_{i+1}$. Using this abstraction, obtaining the optimum for $x_{i+1}$ requires “only” the knowledge of $\alpha_{i}$. Analogously, to calculate $f_{i}$ itself we need to find the optimum of $\alpha{}x_{i}-g_{i}^{*}(\alpha)-\tilde{f}_{i+1}^{*}(\alpha)$ with respect to $\alpha$ which gives us $\alpha_{i}$. This in turn only requires the knowledge of $x_{i}$. To recap, the optimization sub-problem for $x_{i+1}$ solely requires the knowledge of $\alpha_{i}$ whereas the optimization sub-problem for $\alpha_{i}$ solely requires the knowledge of $x_{i}$. Let $\mathsf{x}_{1},\mathsf{a}_{1},\mathsf{x}_{2},\mathsf{a}_{2},\ldots,\mathsf{a}_{n-1},\mathsf{x}_{n}$ be a sequence of locally-admissible pairs $\mathsf{x}_{i},\mathsf{a}_{i}$ and $\mathsf{a}_{i},\mathsf{x}_{i+1}$ where the full sequence is not necessarily globally optimal. The subgradient conditions for pair-restricted optimality define two admissible sets,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\mathsf{x}_{i+1},\mathsf{a}_{i}):\mathsf{a}_{i}\in\partial\tilde{f}_{i+1}(\mathsf{x}_{i+1})\big{\}}$ (5)
$S_{i}$ $=$ $\big{\{}(\mathsf{x}_{i},\mathsf{a}_{i}):\mathsf{a}_{i}\in\partial f_{i}(\mathsf{x}_{i})\big{\}}~{}.$ (6)
@@
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\mathsf{x}_{i+1},\mathsf{a}_{i}):\mathsf{a}_{i}\in\partial\tilde{f}_{i+1}(\mathsf{x}_{i+1})\big{\}}$
@@
$=$ $\big{\{}(\mathsf{x}_{i+1},\mathsf{a}_{i}):\mathsf{a}_{i}-\mathsf{a}_{i+1}\in\partial h_{i+1}(\mathsf{x}_{i+1})\;;\;(\mathsf{x}_{i+1},\mathsf{a}_{i+1})\in S_{i+1}\big{\}}~{}.$
@@
$\mathsf{a}_{i}\in\partial f_{i}(\mathsf{x}_{i})~{}\Leftrightarrow~{}\mathsf{x}_{i}\in\partial g_{i}^{*}(\mathsf{a}_{i})+\partial\tilde{f}_{i+1}^{*}(\mathsf{a}_{i})~{}.$
@@
$=$ $\big{\{}(\mathsf{x}_{i},\mathsf{a}_{i}):\mathsf{x}_{i}\in\partial g_{i}^{*}(\mathsf{a}_{i})+\partial{\tilde{f}_{i+1}}^{*}(\mathsf{a}_{i})\big{\}}$
$=$ $\big{\{}(\mathsf{x}_{i},\mathsf{a}_{i}):\mathsf{x}_{i}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i})\;;\;(\mathsf{x}_{i+1},\mathsf{a}_{i})\in S_{i+\frac{1}{2}}\big{\}}~{}.$
@@
$(\mathsf{x}_{i+1},\mathsf{a}_{i})\in S_{i+\frac{1}{2}}~{}\Leftrightarrow~{}\mathsf{a}_{i}\in\partial\tilde{f}_{i+1}(\mathsf{x}_{i+1})~{}\Leftrightarrow~{}\mathsf{x}_{i+1}\in\partial\tilde{f}^{*}_{i+1}(\mathsf{a}_{i})~{}.$
@@
$\begin{array}[]{lcl}S_{i+\frac{1}{2}}&=&\left\{(\mathsf{x},\mathsf{a}):\mathsf{a}-\mathsf{a}^{\prime}\in\partial h_{i+1}(\mathsf{x})\;;\;(\mathsf{x},\mathsf{a}^{\prime})\in S_{i+1}\right\}\\ S_{i}&=&\left\{(\mathsf{x},\mathsf{a}):\mathsf{x}-\mathsf{x}^{\prime}\in\partial g_{i}^{*}(\mathsf{a})\;;\;(\mathsf{x}^{\prime},\mathsf{a})\in S_{i+\frac{1}{2}}\right\}~{}.\end{array}$
@@
$S_{n}\equiv\big{\{}(\mathsf{x},0):\mathsf{x}\in\text{dom}(h_{n})\big{\}}~{}.$
@@
We henceforth refer to the above general schema as well as the following derived algorithms as Nariat for a shorthand to uNtangled Lariats.

### v2

@@ Abstract
We describe an apparatus for subgradient-following of the optimum of convex problems with variational penalties. In this setting, we receive a sequence $y_{i},\ldots,y_{n}$ and seek a smooth sequence $x_{1},\ldots,x_{n}$. The smooth sequence needs to attain the minimum Bregman divergence to an input sequence with additive variational penalties in the general form of $\sum_{i}{}g_{i}(x_{i+1}-x_{i})$. We derive known algorithms such as the fused lasso and isotonic regression as special cases of our approach. Our approach also facilitates new variational penalties such as non-smooth barrier functions.
We then introduce and analyze new multivariate problems in which ${\mathbf{x}}_{i},\mathbf{y}_{i}\in\mathbb{R}^{d}$ with variational penalties that depend on $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$. The norms we consider are $\ell_{2}$ and $\ell_{\infty}$ which promote group sparsity. We also derive a novel lattice-based procedure for subgradient following of variational penalties characterized through the output of arbitrary convolutional filters. This paradigm yields efficient solvers for high-order filtering problems of temporal sequences in which sparse discrete derivatives such as acceleration and jerk are desirable.
@@ 1 Sub-Introduction
To start, let us examine two seemingly different problems, the fused lasso [25, 10] and isotonic regression [2, 4]. The input to the two procedures is a sequence $\mathbf{y}=y_{1},\ldots,y_{n}$ where the goal is two obtain smoothed sequences, denoted by ${\mathbf{x}}$, that are the solutions of the following optimization problems,
[…]
We can unify the two problems into a single abstract problem of the form,
$\min_{{\mathbf{x}}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}g_{i}(x_{i}-x_{i+1})~{}~{}.$ (1)
Here, both $h_{i}:\mathbb{R}\to\mathbb{R}_{+}$ and $g_{i}:\mathbb{R}\to\mathbb{R}_{+}$ are convex in their single argument. Specific choices for $h_{i}$ and $g_{i}$ yield as special cases the fused-lasso, by setting $g_{i}(x_{i}-x_{i+1})=\lambda_{i}|x_{i}-x_{i+1}|$, and isotonic regression, by defining $g_{i}(x_{i}-x_{i+1})=0$ if $x_{i+1}\geq x_{i}$ and $\infty$ otherwise. Albeit the superficially different form of the variational penalty $g_{i}$, our apparatus facilitates gradient following procedures which differ by a single line of code as shown in Fig. 8. As in the case of isotonic regression, this procedure can facilitate non-smooth variational penalties for $g_{i}$. We fondly refer to the general subgradient-following paradigm as the Untangled Lariat. To illustrate the power of the Untangled Lariat we derive a novel derivative of the generic problem given by Eq. (1) which uses fairly general barrier functions for $g_{i}(\cdot)$.
We next generalize the Untangled Lariat to multivariate settings in which each element of the input sequence is a vector. We thus seek a solution ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}$ where ${\mathbf{x}}_{i}\in\mathbb{R}^{d}$ for all $i\in[n]$. In order to obtain variational sparsity in the multivariate case we replace the absolute value $|x_{i+1}-x_{i}|$ with norm penalties over the differences $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$. To start, we derive a closed-form for the squared $2$-norm, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|^{2}$. We use the closed form solution to build surrogate functions for the $2$-norm itself, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$, and the infinity-norm, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}$. The result is an iterative yet efficient solver for the problem,
@@
For brevity of the derivation of the multivariate setting we confine ourselves to the case where $h_{i}({\mathbf{x}}_{i})=\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}$.
Last, but not least, we introduce a construction from signal processing into our setting by penalizing the convolution of ${\mathbf{x}}$ with a predefined convolution filter $\boldsymbol{\tau}$. Concretely, we provide an efficient procedure for finding the optimum of,
$\min_{\mathbf{x}}\sum_{i}h_{i}(x_{i})+\sum_{i}\big{|}(\boldsymbol{\tau}\ast{\mathbf{x}})_{i+1}\big{|}~{}~{},$
@@
where $\boldsymbol{\tau}\in\mathbb{R}^{k}$ is a predefined yet arbitrary filter of $k-1$ free parameters. We introduce a subgradient-following procedure which operates over a $k$-dimensional lattice. It takes, in the worst case, $\mathcal{O}(n^{k})$ operations to construct the lattice from which the optimal solution is obtained. In applications where $\mathbf{y}$ represents a temporal sequence, the variational penalty of the fused lasso $|x_{i+1}-x_{i}|$ promotes smoothed solutions with regions of zero discrete “velocity”. Analogously, the high-order Untangled Lariat is a filtering and smoothing apparatus for problems in which we seek regions of zero acceleration, zero jerk, and other forms of finite differences.
The initial motivation for this work can be traced back to the influential work on the fused lasso [25, 10] and isotonic regression [2, 4, 29]. The setting of the fused lasso has been revisited numerous times and found various venues for applications. See for instance [12, 13], though any attempt to be exhaustive here is going to do injustice to the plethora of applications that employ the fused lasso. Several authors [8, 16, 18] described and analyzed a linear time algorithm for solving the fused lasso problem which is morally equivalent to Untangled Lariat with the aforementioned specific choices for $h_{i}$ and $g_{i}$. Alternative algorithms for the fused lasso and similar problems, such as the ones described in [3, 7], do not guarantee a runtime linear in $n$ nonetheless demonstrate excellent empirical results.
The paper that is morally most relevant to our work is [18] where the authors follow the derivation in [16] to generalize the algorithm to non-smooth unary functions and to trees. The procedure described in [18] is an instance of a primal-dual message passing algorithm [11]. The authors also consider unary functions subject to a pairwise penalty term which also include the fused lasso penalty as a special case. The lens of the work presented in [18] is probabilistic in nature and does not consider general pairwise penalties.
Several other forms of variational penalties were the focus specific extension of the fused lasso and isotonic regression. For instance, the works presented in [29] considers isotonic regression with $\ell_{1}$ norm for the variational penalty. We readily obtain the same procedure as a special case of our approach. The problem of high-order variational penalties was originally proposed by [17] and its statistical properties were studied by [24, 27]. An algorithmic perspective was provided in [17] which discusses an approximate solution using a primal-dual interior point method. For penalties of the form $\lambda\sum_{i}|x_{i+1}-x_{i}|$, the entire solution path for admissible values of $\lambda$, is provided using a dual path algorithm in [26, 1]. To the best of our knowledge, we give the first efficient analytic solution for a given $\lambda$.
We generalize the aforementioned research papers by providing a single unified algorithmic paradigm for general convex unary functions and a broad class of pairwise terms. Our analysis technique deviates from prior research by employing the Fenchel dual of the primal problem, a view that was advocated in the context of online learning by Shalev-Shwartz [22]. For a thorough study of Fenchel duality and its connection to Lagrange dual see for instance Chapter 3 in [BorweinLewis06]. The incorporation of Fenchel duality leads to a simple and useful generalization to Bregman divergences [6] between ${\mathbf{x}}$ and $\mathbf{y}$ as the objective. This generalization does not bear an additional computation cost. We provide in an appendix Python code that implements subgradient following procedures for most of the settings discussed in this paper.
@@ 2 Abstract Subgradient Following Algorithm
Throughout the paper we assume that $h_{i}$ and $g_{i}$ are convex and lower semi-continuous. The Fenchel dual of a function $f$ is defined as,
$f^{*}(\alpha)=\sup_{x\in\mathrm{dom}(f)}\big{\{}\alpha x-f(x)\big{\}}~{}~{}.$
The domain of the functions we consider is closed, compact, and convex. We can thus replace the supremum above with a maximum. We next utilize Fenchel-Moreau theorem for each $g_{i}$,
$g_{i}(x_{i}-x_{i-1})=\max_{\alpha}\big{\{}\alpha(x_{i}-x_{i-1})-g_{i}^{*}(\alpha)\big{\}}~{}~{},$
in order to evaluate $U$ as follows,
$U$ $=$ $\min_{{\mathbf{x}}}\,\max_{\boldsymbol{\alpha}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}\left[\alpha_{i}(x_{i}-x_{i+1})-g_{i}^{*}(\alpha_{i})\right]~{}~{},$
We confine ourselves to settings for which strong duality holds thus,
$U$ $=$ $\min_{x_{1}}\,\max_{\alpha_{1}}\,\min_{x_{2}}\ldots\max_{\alpha_{n-1}}\,\min_{x_{n}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}\left[\alpha_{i}(x_{i}-x_{i+1})-g_{i}^{*}(\alpha_{i})\right]~{}~{}.$
@@
$f_{i}(x_{i})$$\equiv$ $\max_{\alpha_{i}}\,\ldots\min_{x_{n}}\>\sum_{j=i}^{n-1}\left[h_{j+1}(x_{j+1})+\alpha_{j}(x_{j}-x_{j+1})-g_{j}^{*}(\alpha_{j})\right]~{}~{}.$ (3)
For boundary conditions we let $\forall x:f_{n}(x)=0$. In addition, we define an auxiliary function $f_{i-\frac{1}{2}}(x_{i})=f_{i}(x_{i})+h_{i}(x_{i})$. Using these definitions we have $U=\min_{x_{1}}f_{\frac{1}{2}}(x_{1})$. Next we introduce a summation-free form for $f_{i}$,
@@
$=$ $\max_{\alpha_{i}}\>\alpha_{i}x_{i}-g_{i}^{*}(\alpha_{i})+\min_{x_{i+1}}\left\{f_{i+\frac{1}{2}}(x_{i+1})-\alpha_{i}x_{i+1}\right\}$
$=$ $\max_{\alpha_{i}}\>\alpha_{i}x_{i}-g_{i}^{*}(\alpha_{i})-f_{i+\frac{1}{2}}^{*}(\alpha_{i})~{}~{}.$
Note that to find $f_{i}(x_{i})$ we need to obtain the minimum of $f_{i+\frac{1}{2}}(x_{i+1})-\alpha_{i}x_{i+1}$ with respect to $x_{i+1}$. Using this abstraction, obtaining the optimum for $x_{i+1}$ “only” requires the knowledge of $\alpha_{i}$. Analogously, to calculate $f_{i}$ itself we need to find the optimum of $\alpha{}x_{i}-g_{i}^{*}(\alpha)-f_{i+\frac{1}{2}}^{*}(\alpha)$ with respect to $\alpha$ which gives us $\alpha_{i}$. This in turn only requires the knowledge of $x_{i}$. To recap, the optimization sub-problem for $x_{i+1}$ solely uses the information on $\alpha_{i}$ and analogously the optimization sub-problem for $\alpha_{i}$ only requires information of $x_{i}$. Let $\mathsf{x}_{1},\mathsf{a}_{1},\mathsf{x}_{2},\mathsf{a}_{2},\ldots,\mathsf{a}_{n-1},\mathsf{x}_{n}$ denote a sequence of locally-admissible pairs $\mathsf{x}_{i},\mathsf{a}_{i}$ and $\mathsf{a}_{i},\mathsf{x}_{i+1}$ where the full sequence is not necessarily globally optimal. The subgradient conditions for pair-restricted optimality define two admissible sets,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\mathsf{x}_{i+1},\mathsf{a}_{i}):\mathsf{a}_{i}\in\partial f_{i+\frac{1}{2}}(\mathsf{x}_{i+1})\big{\}}$ (5)
$S_{i}$ $=$ $\big{\{}(\mathsf{x}_{i},\mathsf{a}_{i}):\mathsf{a}_{i}\in\partial f_{i}(\mathsf{x}_{i})\big{\}}~{}~{}.$ (6)
@@
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\mathsf{x}_{i+1},\mathsf{a}_{i}):\mathsf{a}_{i}\in\partial f_{i+\frac{1}{2}}(\mathsf{x}_{i+1})\big{\}}$
@@
$=$ $\big{\{}(\mathsf{x}_{i+1},\mathsf{a}_{i}):\mathsf{a}_{i}-\mathsf{a}_{i+1}\in\partial h_{i+1}(\mathsf{x}_{i+1})\;;\;(\mathsf{x}_{i+1},\mathsf{a}_{i+1})\in S_{i+1}\big{\}}~{}~{}.$
@@
$\mathsf{a}_{i}\in\partial f_{i}(\mathsf{x}_{i})~{}\Leftrightarrow~{}\mathsf{x}_{i}\in\partial g_{i}^{*}(\mathsf{a}_{i})+\partial f_{i+\frac{1}{2}}^{*}(\mathsf{a}_{i})~{}~{}.$
@@
$=$ $\big{\{}(\mathsf{x}_{i},\mathsf{a}_{i}):\mathsf{x}_{i}\in\partial g_{i}^{*}(\mathsf{a}_{i})+\partial{f_{i+\frac{1}{2}}}^{*}(\mathsf{a}_{i})\big{\}}$
$=$ $\big{\{}(\mathsf{x}_{i},\mathsf{a}_{i}):\mathsf{x}_{i}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i})\;;\;(\mathsf{x}_{i+1},\mathsf{a}_{i})\in S_{i+\frac{1}{2}}\big{\}}~{}~{}.$
@@
$(\mathsf{x}_{i+1},\mathsf{a}_{i})\in S_{i+\frac{1}{2}}~{}\Leftrightarrow~{}\mathsf{a}_{i}\in\partial f_{i+\frac{1}{2}}(\mathsf{x}_{i+1})~{}\Leftrightarrow~{}\mathsf{x}_{i+1}\in\partial f^{*}_{i+\frac{1}{2}}(\mathsf{a}_{i})~{}~{}.$
@@
$\begin{array}[]{lcl}S_{i+\frac{1}{2}}&=&\left\{(\mathsf{x},\mathsf{a}):\mathsf{a}-\mathsf{a}^{\prime}\in\partial h_{i+1}(\mathsf{x})\;;\;(\mathsf{x},\mathsf{a}^{\prime})\in S_{i+1}\right\}\\ S_{i}&=&\left\{(\mathsf{x},\mathsf{a}):\mathsf{x}-\mathsf{x}^{\prime}\in\partial g_{i}^{*}(\mathsf{a})\;;\;(\mathsf{x}^{\prime},\mathsf{a})\in S_{i+\frac{1}{2}}\right\}~{}~{}.\end{array}$
@@
$S_{n}\equiv\big{\{}(\mathsf{x},0):\mathsf{x}\in\text{dom}(h_{n})\big{\}}~{}~{}.$